OpenStudy (anonymous):

a woman can do a certain job in 10 hours and her younger daughter can do it in 12 hours. after the woman and her younger daughter have been working for 1 hour, they are join by the older daughter and the three complete the job in 3 more hours. how will it take the older daughter to do the jib by herself. please help i'm really confused.

OpenStudy (anonymous):

do you have the answer?? THis doesn't ha ve a round answer but it is 45/4

OpenStudy (anonymous):

11.25 hrs? how did you get that answer?

OpenStudy (phi):

These are tricky problems. I think of them as rate times time = 1 job where rate is how fast you do a job.

OpenStudy (phi):

so, for example, we have for the first 2 women (job/10 hrs + job/12 hrs) * 1 hr = 22/120 = 11/60 of a job in 1 hour

OpenStudy (phi):

that leaves 1- 11/60, or 49/60 of the job to do. so now we have (job/10 + job/12 + job/x) 3 hrs = (49/60) job where x is the unknown rate.

OpenStudy (phi):

it is now algegra: \[(\frac{1}{10}+\frac{1}{12}+\frac{1}{x})3= \frac{49}{60}\]

OpenStudy (anonymous):

yes - good thinking phi - i find these problems tricky too

OpenStudy (phi):

Can you finish it?

OpenStudy (anonymous):

ok 1/10 + 1/12 + 1/x = 49/180 sorry - gotta go

OpenStudy (anonymous):

18x + 15x + 180 = 48x 49 - 33x = 180 16x = 180 x = 11.25 days for daughter to do job herself

OpenStudy (phi):

Ok, but you do have a few typos. Just to clarify for any viewers, after multiplying both sides by 180x you get 18x+15x+180= 49x 180= 49x-18x-15x 16x=180 x= 11.25